Pop's Cycle Shop sells bicycles and tricycles. The number of bicycles is 1 less than 4 times the number of tricycles. All the bicycles and tricycles together have a total of 174 wheels. How many bicycles are there?
Answer by mananth(16946)

let tricycles be x
bicycles= 4x-1
...
3x+2(4x-1)=174
3x+8x-2=174
11x=174+2
11x=176
x=16 tricycles
bicycles = 4x-1=64-1
= 63 bicycles
...
m.ananth@hotmail.ca